The average hourly wage of plumbers is normally distributed with a population mean of $32.00 and a population standard deviation of $6.00. Calculate the following:
(a) The proportion of plumbers earning between $20 and $24.
(b) The proportion of plumbers earning more than $38.
(c) The proportion of plumbers earning less than $15.
(d) Calculated the 80th percentile.

Answer by edjones(8007)

a)
z=(20-32)/6=-12/6=-2
z=(24-32)/6=-8/6=-4/3
Area under normal curve between -2 and -4/3 is .068 The proportion of plumbers earning between $20 and $24.
.
Ed